The volume of a right circular cone is 5 litres. Calculate the volumes of the two parts into which the cone is divided by a plane parallel to the base, one-third of the way down from the vertex to the base. Give your answers to the nearest ml.
Answer by greenestamps(13198)



The two parts into which the original cone is divided are a small cone similar to the original cone and the frustum that is left.

According to the given information, the height of the small cone is 1/3 of the height of the original cone. By a powerful general principle regarding similar figures, the volume of the small cone is (1/3)^3 = 1/27 of the volume of the original cone.

The volume of the original cone is 5 liters = 5000ml. 1/27 of that volume, rounded to the nearest ml, is 185ml.

So the volume of the small cone is 185ml, and the volume of the frustum is 5000-185 = 4815ml.

ANSWERS: 185ml and 4815ml